# “Niñas Atómicas” (Atomic Girls): An initiative that generates opportunities for young girls in STEM

### v1

@@ IV Results
We present combined results using data taken by the girls during the 2024 and 2025 versions of our workshop, with a total of 50 participants. We show in section IV.1 results on muon flux and proper lifetime, which are two muon properties which can be extracted with the muon counter experiment. In section IV.2, results from a questionnaire that 40 out of the 50 girls responded after finishing the workshop are also presented, related to the girl’s learning perceptions right after completion.
@@ IV.1 Results with the constructed muon detectors
Figure 4 presents the muon data recorded with the muon counter detectors, showing the number of detected muons as a function of time. The dataset from Location A was collected in the Farellones [^10] We thank Renato Galleguillos for taking this data. area near Santiago, at an altitude of 2793 m. The remaining measurements were taken at the San Joaquín campus: Location B at the tallest building (605 m) and Location C in our laboratory (547 m). The dataset from Location A was provided to the students for comparison, while the data from Locations B and C correspond to measurements they carried out at our university. This plot illustrates a representative result that most of the girls were able to produce using their own measurements during the workshop. By analyzing it they can answer their proposed scientific question: Does the number of detected muons depend on the altitude at which the data is collected? The answer is yes: with our muon counter this effect is clearly observed. At higher altitudes, the detector registers more muons. However, when comparing only Locations B and C, the differences in altitude are relatively small, so the effect is less pronounced when compared with Location A.
In addition, the muon lifetime can, in principle, be extracted from the data collected by the students. However, this has proven to be a very challenging task for them, since explaining the phenomenon requires a precise understanding of concepts from special relativity, and not all of them are willing to take on the challenge. Nevertheless, we also include this example to demonstrate the experiment’s full potential. Figure  5 shows the Gaussian fits to data taken at Location A and Location C. The mean of the distributions corresponds to the average number of muons per minute observed at that altitude. At Location A, we detected a mean of $10.1$ muons per minute. This same exercise can be done at several altitudes. Choosing a second measurement at location C we detected a mean of $2.7$ muons per minute. With these two measurements, we can calculate the muon lifetime.
@@
In our example, we assume that the muons travel with a constant velocity of $v_{\mu}=0.9c$, where $c$ denotes the speed of light, and that the time is given by $t=(h_{A}-h_{C})/v_{\mu}$, with $h_{A}$ representing the altitude at Location A and $h_{C}$ the altitude at Location C. For our example, we used $N_{0}=10.1$ and $N(t)=2.7$, corresponding to the mean number of muons observed at altitudes $h_{A}=2793$ m and $h_{C}=547$ m, respectively. From these values, we obtain a muon proper lifetime of:
$\tau=2.7~\text{ns}.$ (3)
@@
Data taking can be further improved. In the 2024 and 2025 editions of the workshop, several measurements were performed by our students over short time intervals of about 20 minutes, in contrast to the data collected at Location A, where the counter operated for approximately one hour. We found that longer acquisition times lead to more stable statistics and more evident Gaussian distributions in the histogram of detected muons per minute, as can bee seen in Figure 5. Therefore, a key lesson learned from this experience is the importance of maximizing the data-taking period of the muon counters whenever possible.
Figure 4: Number of muons versus time at different heights.
Figure 5: Gaussian fits to the data collected at $2793$ m and $547$ m (Locations A and C), where the fitted mean $\mu$ corresponds to the average number of muons detected at that altitude.
@@ IV.2 Workshop evaluation
After each version of the workshop the girl’s responded to a survey, intended for us to understand better their learning experience and to gather feedback on the workshop methodology and its overall dynamics. This also helps us understand if we are meeting the workshop main objectives (see section I). The following results relate to the girls learning perception and workshop appreciation. We present results from the 2024 and 2025 editions of the workshop, which involved a total of 50 participants, of whom 40 completed the survey. Figures 6 and 7 show the level of agreement of the participants to 12 statements. Statements 1–6 align with our first objective of fostering critical thinking and transferable skills. They address the participants’ perceptions of the abilities they believe they acquired through the workshop, as well as how the workshop contributed to their understanding of science and the learning of new concepts. Statements 7–12 are aligned with our second objective of supporting participants in their decision-making process by providing a concrete understanding of what science is and how it is practiced. These items assess their level of agreement regarding how the workshop influenced their perception of science and the work of scientists, their confidence in current or future performance in STEM subjects, and their views on the workshop being exclusively for girls.
@@
In addition to evaluating specific skills, the survey also asked participants to reflect on how the workshop influenced their views on science and on women’s participation in science. As shown in figure 8, a 70% of the girls reported that their perception of science changed “quite a lot” or “very much”, while only 5% indicated no change at all. In figure 9 we see that a 57.5% of respondents indicated that their perceptions concerning the participation of women in science changed “quite a lot” or “very much”, while 20% responded “nothing” . Even though we do not know what are precisely their previous perceptions on these topics prior to the workshop, these responses suggests that most of the participants are, at the very least, questioning their perceptions after participating in the workshop.
@@ V Summary and outlook
We have created an initiative that combines the teaching of particle physics, programming, electronics and scientific methodology through the building of a dedicated particle physics experiment: a muon counter. Our “Niñas Atómicas” workshop has run for two weeks every year since 2022, where high-school girls aged 15 to 16 years old from all over Chile can experience how science works. Our main objectives are to foster transferable scientific skills as well as to support students in their decision-making process when considering a scientific career.
The usage of “hands-on” physics experiments provides an engaging way to introduce students to complex concepts. Our workshop revolves around the muon counter, which is a portable particle detector that the girls use to explore fundamental physics phenomena. We note that (a more complex) pocket-sized muon detector design is described in Axani:2018qzs, which could be also used in initiatives of this kind. Other initiatives that include the use of interactive physics apparatus at schools suggest an increased interest in the pursuit of a STEM career Lubrica2017. Teaching particle physics at the high-school level is also a valuable experience for diversifying students backgrounds in physics before entering the university RogerBarlow_1992, and have had a long-term impact worldwide, for instance, with the development of the Particle Physics Masterclasses Masterclasses; Bilow:2022rs; 2024arXiv240117103T. Moreover, initiatives of this kind can aid bring young people awareness on High Energy Physics (HEP) in Latinamerica Assamagan:2023eac.
@@
Other aspect to consider is that the girls must have access to a computer with an internet connection. We do require this at the time of application, so their effective participation relies on them to be able to access a computer (not cellphones or tablets) during the remote sessions at their homes, high schools or local communities. An additional challenge for us was a limited amount of reliable resources on particle physics in Spanish. This meant to dedicate particular efforts in making our lessons and material as self-contained as possible. As most of the girls that participated do not speak well English, and most science is done and communicated in English, we aim to contribute to narrow this gap in resources with time. Our detector assembly manual and particle physics handbook can be found in AtomicasGithub.
@@
None of the above challenges were limiting factors in the creation of this initiative, which we hope remains stable in time. Our overarching goal is to cultivate scientific thinking among young minds in Chile, enabling them to become ambassadors at their local communities as representatives of our “Niñas Atómicas” program. The final scientific reports on their work at the workshop has become a valuable resource, as the girls themselves can share them and present them at their high schools and local scientific fairs. We hope to keep encouraging students to ask questions, develop the tools and skills to be able to answer them, and more importantly, to inspire them to remain curious about the universe we live in.
@@ Acknowledgments
We thank Renato Galleguillos and the Center for Theoretical and Experimental Particle Physics (CTEPP) at Universidad Andrés Bello (UNAB) for the design of the PCB, which is essential for this muon counter experiment. We thank Florencia Díaz, our electronics teacher and Suyay Huichacura, Melanie Martínez, Devika Muhki and Mariel Poduje, our senior tutors of the “Niñas Atómicas” program. We also thank Roberto Pinto and the Institute of Physics at Pontificia Universidad Católica de Chile for providing laboratory space and financial support in the developing of the 2024 and 2025 versions of the workshop. We also acknowledge financial support from ANID FONDECYT grant No. 1240721 and ANID – Millennium Science Initiative Program ICN2019_044, as well as the Vice-Rectorate of Research at Pontificia Universidad Católica de Chile.
@@
This study was approved by the Vice-Rectorate of Research at Pontificia Universidad Católica de Chile and by the program “Proyección al Medio Externo” (PME) by ANID Millennium Science Initiative Program. All parents, guardians or next of kin provided written informed consent for the minors to participate in this study. All adult participants provided written informed consent to participate in this study and we have consent for the student’s results to be published.

### v2

@@ IV Results
We present combined results using data taken by the girls during the 2024 and 2025 versions of our workshop, with a total of 50 participants. We show in section IV.1 results on muon flux and proper lifetime, which are two muon properties which can be extracted with the muon detector. In section IV.2, results from a questionnaire that 40 out of the 50 girls responded after finishing the workshop are also presented, related to the girl’s learning perceptions right after completion.
@@ IV.1 Results with the constructed muon detectors
Figure 4 presents the muon data recorded with the muon detectors, showing the number of detected muons as a function of time. The dataset from Location A was collected in the Yerba Loca park in Farellones near Santiago, at an altitude of 1850 m. The remaining measurements were taken at the San Joaquín campus: Location B at the tallest building (605 m) and Location C in our laboratory (547 m). The dataset from Location A is provided for comparison, while the data from Locations B and C correspond to measurements the girls carried out at our university. This plot illustrates a representative result that most of the girls were able to produce using their own measurements during the workshop. By analyzing it they can answer their proposed scientific question: Does the number of detected muons depend on the altitude at which the data is collected? The answer is yes: with our muon detector this effect is clearly observed. At higher altitudes, the detector registers more muons. However, when comparing only Locations B and C, the differences in altitude are relatively small, so the effect is less pronounced when compared with Location A.
In addition, the muon lifetime can, in principle, be extracted from the data collected by the students. However, this has proven to be a very challenging task for them, since explaining the phenomenon requires a precise understanding of concepts from special relativity (which are very unfamiliar to most students). Nevertheless, we also include this example to demonstrate the experiment’s full potential. Figure  5 shows the Gaussian fits to data taken at Location A and Location C. The mean of the distributions corresponds to the average number of muons per minute observed at that altitude. At Location A, we detected a mean of $9.1$ muons per minute. This same exercise can be done at several altitudes. Choosing a second measurement at location C we detected a mean of $2.7$ muons per minute. With these two measurements, we can calculate the muon lifetime.
@@
In our example, we assume that the muons travel with a constant velocity of $v_{\mu}=0.9c$, where $c$ denotes the speed of light, and that the time is given by $t=(h_{A}-h_{C})/v_{\mu}$, with $h_{A}$ representing the altitude at Location A and $h_{C}$ the altitude at Location C. For our example, we used $N_{0}=9.1$ and $N(t)=2.7$, corresponding to the mean number of muons per minute observed at altitudes $h_{A}=1850$ m and $h_{C}=547$ m, respectively. From these values, we obtain a muon proper lifetime of:
$\tau=1.7~\text{ns}.$ (3)
@@
Data taking can be further improved. In the 2024 and 2025 editions of the workshop, several measurements were performed by our students over short time intervals of about 20 minutes, in contrast to the data collected at Location A, where the counter operated for approximately one hour. We found that longer acquisition times lead to more stable statistics and more evident Gaussian distributions in the histogram of detected muons per minute, as can bee seen in Figure 5. Therefore, a key lesson learned from this experience is the importance of maximizing the data-taking period of the muon detector whenever possible.
Figure 4: Number of muons versus time at different heights. Location A: Yerba Loca park in Farellones near Santiago, at an altitude of 1850 m. Location B: tallest building at San Joaquín campus (605 m). Location C: physics laboratory at San Joaquín campus (547 m).
Figure 5: Gaussian fits to the data collected at $1850$ m and $547$ m (Locations A and C), where the fitted mean $\mu$ corresponds to the average number of muons per minute detected at that altitude.
@@ IV.2 Workshop evaluation
After each version of the workshop the girls responded to a survey, intended for us to understand better their learning experience and to gather feedback on the workshop methodology and its overall dynamics. This also helps us understand if we are meeting the workshop main objectives (see section I). The following results relate to the girls learning perception and workshop appreciation. We present results from the 2024 and 2025 editions of the workshop, which involved a total of 50 participants, of whom 40 completed the survey. Figures 6 and 7 show the level of agreement of the participants to 12 statements. Statements 1–6 align with our first objective of fostering critical thinking and transferable skills. They address the participants’ perceptions of the abilities they believe they acquired through the workshop, as well as how the workshop contributed to their understanding of science and the learning of new concepts. Statements 7–12 are aligned with our second objective of supporting participants in their decision-making process by providing a concrete understanding of what science is and how it is practiced. These items assess their level of agreement regarding how the workshop influenced their perception of science and the work of scientists, their confidence in current or future performance in STEM subjects, and their views on the workshop being exclusively for girls.
@@
In addition to evaluating specific skills, the survey also asked participants to reflect on how the workshop influenced their views on science and on women’s participation in science. As shown in figure 8, a 70% of the girls reported that their perception of science changed “quite a lot” or “very much”, while only 5% indicated no change at all. In figure 9 we see that a 57.5% of respondents indicated that their perceptions concerning the participation of women in science changed “quite a lot” or “very much”, while 20% responded “nothing” . Even though we do not present results at this time on what are precisely their previous perceptions prior to the workshop (another work is in progress in EstudioImpacto), these responses suggests that most of the participants are, at the very least, questioning their perceptions after participating in the workshop.
@@ V Summary and outlook
We have created an initiative that combines the teaching of particle physics, programming, electronics and scientific methodology through the building of a dedicated particle physics experiment: a muon detector that counts muons. Our “Niñas Atómicas” workshop has run for two weeks every year since 2022, where high-school girls aged 15 to 16 years old from all over Chile can experience how science works. Our main objectives are to foster transferable scientific skills as well as to support students in their decision-making process when considering a scientific career.
The usage of “hands-on” physics experiments provides an engaging way to introduce students to complex concepts. Our workshop revolves around building a muon detector, which is a portable particle detector that the girls use to explore fundamental physics phenomena. We note that a simple experimental setup for cosmic muon lifetime measurements is also described in Bosnar:2018ngm. Other similar experiments are described in Coan_2006; Singh2015MuonDetection; Barazandeh2016MuonTeaching; Axani_2017. A more complex pocket-sized muon detector design is also described in Axani:2018qzs. All these muon detector setups could be used in initiatives of this kind. Other initiatives that include the use of interactive physics apparatus at schools suggest an increased interest in the pursuit of a STEM career Lubrica2017. Teaching particle physics at the high-school level is also a valuable experience for diversifying students backgrounds in physics before entering the university RogerBarlow_1992, and have had a long-term impact worldwide, for instance, with the development of the Particle Physics Masterclasses Masterclasses; Bilow:2022rs; 2024arXiv240117103T. Moreover, initiatives of this kind can aid bring young people awareness on High Energy Physics (HEP) in Latinamerica Assamagan:2023eac.
@@
Other aspect to consider is that the girls must have access to a computer with an internet connection. We do require this at the time of application, so their effective participation relies on them to be able to access a computer (not cellphones or tablets) during the remote sessions at their homes, high schools or local communities. An additional challenge for us was a limited amount of reliable resources on particle physics in Spanish. This meant to dedicate particular efforts in making our lessons and material as self-contained as possible. As most of the girls that participated do not speak well English, and most science is done and communicated in English, we aim to contribute to narrow this gap in resources with time. A detector assembly manual and particle physics handbook created by our tutors can be found in AtomicasGithub.
@@
None of the above challenges were limiting factors in the creation of this initiative, which we hope remains stable in time. Our overarching goal is to cultivate scientific thinking among young minds in Chile, enabling them to become ambassadors at their local communities as representatives of our “Niñas Atómicas” program. The final scientific reports on their work at the workshop have become a valuable resource, as the girls themselves can share them and present them at their high schools and local scientific fairs. We hope to keep encouraging students to ask questions, develop the tools and skills to be able to answer them, and more importantly, to inspire them to remain curious about the universe we live in.
@@ Acknowledgments
We thank Renato Galleguillos and the Center for Theoretical and Experimental Particle Physics (CTEPP) at Universidad Andrés Bello (UNAB) for the design of the muon detector and the PCB, which is essential for this muon experiment. We thank Florencia Díaz (M.Sc. student in Physics, UC), our electronics teacher and Suyay Huichacura (M.Sc. student in Medical Physics, UC), Melanie M. Villarreal (Ph.D. (c) in Physics, UC), Devika Mukhi-Nilo (M.Sc. student in Astrophysics, UC) and Mariel Poduje (Ph.D. (c) in Physics, UC), our senior tutors of the “Niñas Atómicas” program for guiding the students and aiding them in the data taking, as well as for many useful discussions concerning the workshop over the past years. In particular, we thank Mariel Poduje and Devika Mukhi-Nilo for the Spanish version of the manual. We are grateful to Gabriel Torrealba, Florencia Díaz and Devika Mukhi-Nilo for useful comments on this manuscript. We thank Ignacio García for taking updated photographs of the experiment and his review of the detector manual and detailed instructions in English. We also acknowledge our administrative and logistics support team at SAPHIR Millennium Institute. We thank Roberto Pinto for his constant support of our initiative and the Institute of Physics at the Faculty of Physics of Pontificia Universidad Católica de Chile for providing laboratory space and financial support in the developing of the 2024 and 2025 versions of the workshop. We also acknowledge financial support from ANID FONDECYT grant No. 1240721 and ANID – Millennium Science Initiative Program ICN2019_044, as well as the Vice-Rectorate of Research at Pontificia Universidad Católica de Chile.
@@
This study was approved by the Vice-Rectorate of Research at Pontificia Universidad Católica de Chile and by the program “Proyección al Medio Externo” (PME) by ANID Millennium Science Initiative Program. All parents, guardians or next of kin provided written informed consent for the minors to participate in this study and we have consent for the student’s results to be published.